### v1

# Generating quantum non-local entanglement with mechanical rotations

Abstract
Recent experiments have searched for evidence of the impact of non-inertial motion on the entanglement of particles. The success of these endeavours has been hindered by the fact that such tests were performed within spatial scales that were only "local" when compared to the spatial scales over which the non-inertial motion was taking place. We propose a Sagnac-like interferometer that, by challenging such bottlenecks, is able to achieve entangled states through a mechanism induced by the mechanical rotation of a photonic interferometer. The resulting states violate the Bell-Clauser-Horne-Shimony-Holt (CHSH) inequality all the way up to the Tsirelson bound, thus signaling strong quantum nonlocality. Our results demonstrate that mechanical rotation can be thought of as resource for controlling quantum non-locality with implications also for recent proposals for experiments that can probe the quantum nature of curved spacetimes and non-inertial motion.
I Introduction
The seminal work of J. S. Bell allowed to infer the inherent incompatibility of quantum mechanics with the (classically acceptable) assumption of local realism posed by Einstein, Podolsky, and Rosen [1, 2]. The falsification of a Bell inequality, which would be fully satisfied by any local realistic theory, has been reported in countless experiments [3, 4, 5, 6, 7, 8, 9, 10, 9, 11], and recognised with the 2022 Nobel prize in Physics.
Independently, questions about relativity led Sagnac to establish a now widespread method for measuring rotational motion using optical interferometry [12, 13]. Two counter-propagating signals acquire a phase difference proportional to the angular frequency of rotation [14, 15, 16, 17]. This insight led to the development of the ring laser [18] and fiber-optical gyroscopes  [19, 20], with the current state-of-the-art achieving sub-shot-noise sensitivities [21]. The Sagnac effect has been shown to induce interference at the level of quantum systems, with experimental implementations in matter-wave interferometry [17] and single-photon platforms [22].
More recently, photonic technologies have enabled the exploration of rotation-induced quantum phenomena with two-photon experiments. Polarization-entangled photon pairs were shown to be robust against a 30g acceleration achieved on a rotating centrifuge [23]. Using a Hong-Ou-Mandel interferometer on a rotating platform it was found that low frequency mechanical rotations affect bunching statistics [24]. Super-resolution and Sagnac phase sensitivity beyond the shot-noise limit was achieved in [25] using path-entangled NOON states, and milli-radian phase resolution was achieved in [26], allowing the measurement of the Earth’s angular frequency of $\sim 10\mu\text{Hz}$. Furthermore, it was suggested that photonic entanglement can be revealed or concealed using non-inertial motion accessible to current experiments [27]. Using a Hong-Ou-Mandel interferometer with nested arms it was demonstrated that photonic behavior can change from bunching to antibunching (i.e., from bosonic to fermionic) solely due to mechanical rotations [28]. Moreover, it was shown that rotational motion can change the phase of polarization entangled states enabling transitioning between pairs of Bell states [29].
A step further is to demonstrate the actual generation of entanglement using non-inertial motion. The approach proposed in Ref. [30] made use of a multi-path Sagnac interferometer to achieve a maximally entangled path-polarization state of a single-photon. Such state would be suitable for quantum non-contextuality tests aimed at ascertaining whether observables can be assigned preexisting values prior to measurements [31]. In principle, the generated entanglement could be also transferred to two spatially separated physical systems [32], but the experimental demonstration of such procedures remains experimentally challenging [33, 34]. It is thus not immediately obvious whether the single-photon scheme proposed in  [30] would allow to unambiguously demonstrate the generation of genuine quantum (nonlocal) entanglement as opposed to local entanglement.
In a different context, recent theoretical studies in quantum gravity [35, 36] have proposed schemes where the generation of two-particle nonlocal entanglement can be used to witness the quantumness of the gravitational interaction mediator. These proposals fit within a more general framework of studies where researchers pursue methodologies to test the quantumness or nonclassicality of the involved parties [37], regardless of the type of interaction (see also results in optomechanics and biophotonics [38, 39]).
Inspired by these two-party schemes, in this work we show that it is possible to use mechanical rotations to generate and control nonlocal quantum entanglement in an experimental regime that is fully accessible to current photonic technology. We propose a simple scheme consisting of a single Sagnac fibre loop, linear optics elements, photon-pair sources, and a Bell-test detection setup, all placed on a rotating platform. We find that an initially separable two-photon state can be transformed into a maximally entangled state of polarization that violates significantly the Bell-Clauser-Horne-Shimony-Holt (CHSH) inequality [40] as a function of the angular frequency of rotation, obtaining a simple formula for the frequency required to saturate Tsirelson’s bound [41].
Figure 1: (a) Photonic setting for the rotation-controlled generation of quantum non-local states of polarization. Two photons are initially prepared in the separable state $|HV\rangle$ and injected into the setup. Beam-splitters BS1 and BS2 send the photons into a Sagnac loop, and then redirect them towards two individual detection stages. As the photons entering the Sagnac loop through Beam-splitter BS3 are initially prepared in orthogonal polarization states, they do not interact at any point via electromagnetic couplings. We measure the polarization of the photons with a standard Bell detection scheme. One photon is measured using the setup at the top (managed by Alice) and one photon is measured with the setup on the right (managed by Bob). (b) Theoretical prediction of the violation of the Bell-CHSH inequality (blue) and the probability of detection (orange) [cf. Eq. (11) and (13), respectively]. For concreteness, we have set the photon wavelength to $\lambda=1\,\text{$\mu$m}$ ($\omega=2\pi c/\lambda$ ) and the interferometric area to $\sim 7.8\,\text{m}^{2}$ (e.g., $10$ loops of fiber with radius $r=0.5\,\text{m}$ with a total length of $\sim 31.5\,\text{m}$). The loss of a photon pair occurs when the beam-splitters do not direct the photons to Alice and Bob, thus lowering the probability of detection. We predict that a violation ($|{S}|>2$) occurs periodically with the rotation frequency as stated by Eq. (12). The Tsirelson’s bound is first achieved for $\Omega_{\text{Bell}}\sim 0.4\,\text{Hz}$.
II Proposed scheme.
We consider the setup shown in Fig. 1 (a). It consists of two separate lasers, each pumping a non-linear crystal that, through Type I spontaneous parametric down-conversion, generates photon pairs of the same polarization of which we keep just one photon. The signals thus generated are first directed by beam-splitters (BSs) into a Sagnac loop, and then towards two spatially separated Bell-detection apparatuses. We assume that two independent photons are initially prepared in the separable state:
$|\psi_{\text{i}}\rangle=\hat{a}_{H}^{\dagger}\hat{b}_{V}^{\dagger}|0\rangle\equiv|H\,V\rangle,$ (1)
where $\hat{a}_{H}$ ($\hat{b}_{V}$) denotes horizontal H (vertical V) polarization mode. We assume that the two photons propagate past beamsplitters BS1 and BS2 retaining the form of the state in Eq. (1). As the two photons enter the Sagnac loop through BS3, the state changes according to a beam-splitter transformation into
$|\psi_{\text{i}}\rangle\rightarrow\ket{\psi_{1}}=\frac{1}{2}(\hat{a}_{H}^{\dagger}+i\hat{b}_{H}^{\dagger})(i\hat{a}_{V}^{\dagger}+\hat{b}_{V}^{\dagger})|0\rangle,$ (2)
where $\hat{a}$ ($\hat{b}$) denote the co-rotating (counter-rotating) mode. The effect of mechanical rotation is to introduce Sagnac phases with a sign depending on the sense of motion of the particular mode [27]. From Eq. (2) we thus find
|$\ket{\psi_{1}}{\to}\ket{\psi_{2}}{=}\frac{1}{2}\left(e^{i\frac{\phi}{2}}\hat{a}_{H}^{\dagger}{+}ie^{-i\frac{\phi}{2}}\hat{b}_{H}^{\dagger}\right)\left(ie^{i\frac{\phi}{2}}\hat{a}_{V}^{\dagger}{+}e^{-i\frac{\phi}{2}}\hat{b}_{V}^{\dagger}\right)\!|0\rangle,$ (3)
where the phase factors have been introduced to account for the relative phase acquired by the counter-propagating modes. Irrespective of the medium, shape of the interferometer or the location of the center of rotation, the Sagnac phase is given by [42]
$\phi=\frac{4A\omega\Omega}{c^{2}},$ (4)
where $A$ is the interferometer area, $c$ is the speed of light in vacuum, $\omega=2\pi c/\lambda$ is the angular frequency of the photons ($\lambda$ is the photon wavelength), $\Omega=2\pi f$, and $f$ is the mechanical frequency of the rotating platform. Moreover, any phase affecting differently the two polarizations would simply factor out of Eq. (3). Similarly, if there is some random noise affecting the co-rotating path, then the same noise will also affect the counter-rotating path, again factoring out of Eq. (3). Hence, classical phase delays arising through experimental imperfections will not change the final result. To generate differential phases depending on the direction of rotation the only plausible mechanism is the Sagnac effect.
As the photons exit the central loop through BS3, we apply the inverse beam-splitter transformation to Eq. (3), giving us
$\frac{1}{4}$ $\left[\left(e^{i\phi/2}(-i\hat{a}_{H}^{\dagger}+\hat{b}_{H}^{\dagger})+ie^{-i\phi/2}(\hat{a}_{H}^{\dagger}-i\hat{b}_{H}^{\dagger})\right)\right.$ (5)
×$\times\left.\left(ie^{i\phi/2}(-i\hat{a}_{V}^{\dagger}+\hat{b}_{V}^{\dagger})+e^{-i\phi/2}(\hat{a}_{V}^{\dagger}-i\hat{b}_{V}^{\dagger}\right)\right]|0\rangle,$
which, after rearranging, gives
$\left(\frac{1}{2}\left[(\text{cos}\phi-1)\hat{a}_{H}^{\dagger}\hat{b}_{V}^{\dagger}+(\text{cos}\phi+1)\hat{b}_{H}^{\dagger}\hat{a}_{V}^{\dagger}\right]\right.$
$\left.+\frac{\text{sin}\phi}{2}\left[\hat{a}_{H}^{\dagger}\hat{a}_{H}^{\dagger}-\hat{b}_{V}^{\dagger}\hat{b}_{V}^{\dagger}\right]\right)$ $|0\rangle.$ (6)
The first line of Eq. (6) represents the case when each pair of detectors detects one photon, i.e., Alice detects the mode $a$ and Bob detects mode $b$ or vice versa, while the second line represents the case when both photons arrive at the same pair of detectors, i.e., either both to Alice or both to Bob. As we want to compute quantum correlations between Alice and Bob we consider the state conditional on the postselection of the events that provide coincidences at the detectors, which reads
$|\psi_{\text{f}}\rangle$ $=\frac{\text{cos}(\phi)+1}{\sqrt{2(\text{cos}(\phi)^{2}+1)}}|H\,V\rangle+\frac{\text{cos}(\phi)-1}{\sqrt{2(\text{cos}(\phi)^{2}+1)}}|V\,H\rangle,$ (7)
where we have used $\hat{a}_{H}^{\dagger}\hat{b}_{V}^{\dagger}|0\rangle\equiv|H\,V\rangle$ and $\hat{a}_{V}^{\dagger}\hat{b}_{H}^{\dagger}|0\rangle\equiv|V\,H\rangle$, and included the overall state normalization for completeness.
We first note that for $\phi=0$ (corresponding to the case without mechanical rotation) we always remain in the initial state $|H\,V\rangle$, which is separable. More generally, we note that for $\phi=\pi\,k$ ($k\in\mathcal{Z})$ we remain either in the initial state $|H\,V\rangle$ ( $k$ even) or transform into the flipped polarization state $|V\,H\rangle$ ( $k$ odd). However, for any other value of $\phi$ we find that Eq. (7) will be in an entangled state. In particular, for $\phi=\pi/2+\pi\,k$ ($k\in\mathcal{Z})$ Eq. (7) transforms into the maximally entangled Bell state
$|\psi_{\text{f}}\rangle=\frac{1}{\sqrt{2}}\left(|H\,V\rangle-|V\,H\rangle\right),$ (8)
which is usually denoted as the $|\Psi^{-}\rangle$ state.
Our results depend critically on two steps: (i) on the post-selection step from Eq. (6) to (7), where we have discarded the photons that are not going to Alice and Bob, and (ii) on a non-zero rotationally induced phase $\phi\neq 0$, which arises only for non-zero frequencies of rotation $\Omega\neq 0$. Importantly, the post-selection step is not enough to induce entanglement in the absence of mechanical rotation as discussed above for the case $\phi=0$.
III Violation of the Bell-chsh inequality
To quantify the degree of generated non-locality we can compute the Bell-CHSH function $S$ [40]
$S=E(\bm{a},\bm{b})-E(\bm{a},\bm{b}^{\prime})+E(\bm{a}^{\prime},\bm{b})+E(\bm{a}^{\prime},\bm{b}^{\prime}),$ (9)
where
$E(\bm{a},\bm{b})=\langle\psi_{\text{f}}|(\bm{a}\cdot\bm{\sigma})\otimes(\bm{b}\cdot\bm{\sigma})|\psi_{\text{f}}\rangle$ (10)
denote quantum correlation functions. Here, $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ is the vector of the Pauli matrices and $\bm{a}$,$\bm{a}^{\prime}$,$\bm{b}$,$\bm{b}^{\prime}$ are vectors whose direction determines the value of $S$. In order to achieve the highest violation of the Bell test, we choose the vectors $\bm{a}=(1,0,0)$, $\bm{a}^{\prime}=(0,1,0)$, $\bm{b}=(1,1,0)/\sqrt{2}$, $\bm{b}^{\prime}=(-1,1,0)/\sqrt{2}$, thus obtaining
$S=4\sqrt{2}\frac{\sin^{2}(\phi)}{3+\cos(2\phi)}.$ (11)
Local realism enforces the Bell-CHSH inequality $|S|\leq 2$. However, for suitable choices of $\phi$, Eq. (11) allows to violate such constraint. In particular, as stated in Eq. (8), by setting $\phi=\pi/2+k\pi\leavevmode\nobreak\ (k\in\mathcal{Z})$, the state in Eq. (7) reduces to the Bell state $|\Psi^{-}\rangle$ and we achieve the notorious Tsirelson’s bound $|S|=2\sqrt{2}$ [41]. This occurs when the mechanical frequency $\Omega$ takes the values
$\Omega_{\text{Bell}}\equiv\frac{\pi c^{2}}{8A\omega}(2k+1),\qquad(k\in\mathcal{Z}).$ (12)
Here, $k<0$ ($k>0$) would correspond to an (anti-) clockwise sense of rotation. Needless to say, the postselection process required to get Eq. (7) entails that only coincidence events should be considered in order to construct $S$. These occur with a overall detection probability
$P=\frac{1+\text{cos}(\phi)^{2}}{32}.$ (13)
In Fig. 1(b), we plot the Bell-CHSH function and such probability of detection for a set of values of the relevant physical parameters that are well within reach of existing photonic technology. We remark that photon losses would only lower the probability of detection, without affecting the quality of the resulting state achieved through our scheme, which is thus robust against the most relevant source of imperfections in our chosen experimental platform.
IV Discussion
We have proposed a method for the controlled generation of non-locality using mechanical rotation that achieves the map
$|H\,V\rangle\rightarrow(\text{cos}(\phi)+1)|H\,V\rangle+(\text{cos}(\phi)-1)|V\,H\rangle,$ (14)
where we have omitted the normalization for brevity (see Eqs. (1) and (7)), and we recall that $\phi\propto\Omega$ is the Sagnac phase proportional to the angular frequency of the mechanical rotation $\Omega$. By controlling the value of $\Omega$ we can thus prepare separable or nonlocally entangled final states.
The map in Eq. (14) satisfies a number of desiderata: (i) For $\phi=0$ the transformation reduces to the identity map. In other words, without mechanical rotation, the state remains invariant (and classical). (ii) For $\phi=\pi\,k$ ($k\in{\mathbb{Z}})$ the transformation is either the identity operation (even $k$) or induces a polarization flip (odd $k$). The latter case shows that mechanical rotation can be used to swap the polarization state of photon pairs. (iii) For $\phi=(2k+1)\pi/2$ ($k\in{\mathbb{Z}})$ we generate the Bell state $|\Psi^{-}\rangle$. The generated state is expected to induce a maximal violation of the CHSH inequality given by the Tsirelson’s bound [41]. Importantly, the state $|\Psi^{-}\rangle$ arises only when we tune the mechanical frequency of rotation to the value $\Omega_{\text{Bell}}$ given in Eq. (12).
The scheme is also robust against imperfections and noise due to the inherent protection characteristic of the Sagnac loop. Suppose some unwanted phases would be accumulating depending on the polarization $H$, $V$; this would contribute only to a global phase in Eq. (3), but no measurable differential phase would be generated. Similarly, any other random phase affecting the co-rotating path will automatically affect also the counter-rotating path, thus factoring out without affecting the final state. As shown in Fig. 1(b) the experimental parameters required to test the maximum violation of the CHSH inequality can be achieved with current photonic technologies by adaptation of previous experimental schemes [23, 28, 25, 26, 29, 24]. As such we do not expect any new fundamental or technical issue in the implementation of this proposal.
A further benefit of the proposed scheme is also that it does not rely on specific models, but rather on the well established CHSH test. Naively, one would think that the post-selection step from Eq. (6) to Eq. (7) might be responsible for generating non-locality. However, this is not the case as without rotation, Eq. (7) reduces to a separable state. Hence non-zero mechanical rotation is a critical factor for generating non-locality in this setup, i.e., we can legitimately speak of rotationally induced nonlocality. Furthermore, it has been shown that in case of high-efficiency detectors, the violation of the CHSH inequality provides a measurement of nonlocality also in the frame of post-selection [43]. Other types of inequalities can be considered in the experiments to account for inefficient quantum detectors [43].
The question of how to interpret the experiment is of course nonetheless interesting. In this work we have provided a simple yet very effective and powerful theoretical interpretation only relying on the Sagnac phase. While here we have not shown this, the Sagnac phase is of intrinsic relativistic origin. Evidence of such nature stems from Eq. (4), which depends on the speed of light in vacuum and not on that of photons in a medium, suggesting that its origin is related to the spacetime metric (we refer the interested reader to the reviews [14, 15, 16, 17]). However, more formal interpretations within quantum theory in curved space [27], broader quantum field theoretic framework [44], or a general relativistic context [45, 46, 47, 48] are also possible. The possibility to further such thoughts and interpret the spacetime metric as in a superposition, along the lines of Ref. [30], thus reaching out to the domain of quantum reference [49, 50] frames, will be the topic of further investigations.
Acknowledgements
MT acknowledges funding by the Leverhulme Trust (RPG-2020-197). MP acknowledges the support by the Horizon Europe EIC Pathfinder project QuCoM (Grant Agreement No. 101046973), the Leverhulme Trust Research Project Grant UltraQuTe (grant RGP-2018-266), the Royal Society Wolfson Fellowship (RSWF/R3/183013), the UK EPSRC (EP/T028424/1), and the Department for the Economy Northern Ireland under the US-Ireland R&D Partnership Programme. DF acknowledges support from the Royal Academy of Engineering and the UK EPSRC (EP/W007444/1).